Grade 11Mechanics

T=px^2+qx+r p q r are constants. What is the relationship between velocity and acceleration

Animesh Pandey
10 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
T=px^2+qx+r
velocity,
differnetiate w.r.t time we get,
dt = (2px + q)dx
dx/dt = 1/(2px + q)
v = 1/(2px +q)
a = vdv/dx
substitte  and get value hence.